# Relevance-driven Decision Making for Safer and More Efficient Human Robot Collaboration

### v1

@@ V-B Real-time and Safer Decision Making
Through all the test cases, our methodology of relevance determination robustly and correctly predicts the human’s objective and the set of relevant elements for making cereals as bowl, spoon, milk, and cereal. Our human robot task allocation module robustly solves the optimization problem in real-time to generate the tasks for the robot. Thus, we focus on the evaluation of the virtual obstacle and collision avoidance. The test results of our real-time and safe decision making methodology are shown in Table II.
@@
TABLE II: Comparison of Collision Rate. With relevance and future trajectory of human motion considered, the safety of HRC can be dramatically improved.
@@
VI Conclusions and Future Works
Relevance is a novel concept inspired by the human cognitive abilities to evaluate the pertinence of objects based on the human’s objective or other factors. In this paper, we developed a novel two-loop framework that robustly, efficiently, and accurately quantifies relevance and applies relevance to improve the effectiveness and safety of HRC. This framework integrates an asynchronous loop to leverage world knowledge from an LLM and quantify relevance, as well as a real-time loop to execute scene understanding, human intention prediction, and decision-making. Moreover, we proposed and developed a decision making methodology based on relevance, integrating human robot task allocation and real-time motion generation. In motion generation, we developed a methodology to construct a virtual obstacle and formulate the associated repulsive force. Simulations and experiments verify that our framework performs well for relevance quantification, with an objective prediction accuracy of 0.90 and a relevance prediction accuracy of 0.96. Simulations further verify our novel motion generation methodology dramatically decreases the cases with a collision by 63.76% and the frames with a collision by 44.74%. The robot is comprehensively informed about how to best assist the human with relevance, and our decision making module generates safe actions to achieve the assistance.

### v2

@@ V-B Real-time and Safer Decision Making
Through all the test cases, our methodology of relevance determination robustly and correctly predicts the human’s objective and the set of relevant elements for making cereals to be bowl, spoon, milk, and cereal. Our human robot task allocation module robustly solves the optimization problem in real-time to generate the tasks for the robot. Thus, we focus on the evaluation of the virtual obstacle and collision avoidance. The test results of our real-time and safe decision making methodology are shown in Table II.
@@
TABLE II: Comparison of Collision Rate. With relevance and RAPF, the safety of HRC can be dramatically improved.
@@
VI Conclusions
Relevance is a novel concept inspired by human cognitive abilities, enabling adaptive dimensionality reduction of a scene based on objectives, context, or other influencing factors. In this paper, we developed a novel two-loop framework that robustly, efficiently, and accurately quantifies relevance and applies relevance to improve the effectiveness and safety of HRC. This framework integrates an asynchronous loop to leverage world knowledge from an AI toolkit and quantify relevance, as well as a real-time loop to execute scene understanding, human intention prediction, and decision-making. Moreover, we proposed and developed a decision making methodology based on relevance, integrating human robot task allocation and real-time motion generation. In motion generation, we developed a methodology to construct a virtual obstacle and formulate the associated repulsive force. Simulations and experiments verify that our framework performs well for relevance quantification, with an objective prediction accuracy of 0.90 and a relevance prediction accuracy of 0.96. Simulations further verify our novel motion generation methodology dramatically decreases the cases with a collision by 63.76% and the frames with a collision by 44.74%. The robot is comprehensively informed about how to best assist the human with relevance, and our decision making module generates safe actions to achieve the assistance, leading robotics toward artificial general intelligence.